OpenStudy (anonymous):

how do i express f(x) = x 2 + 4x − 6 in standard form ?

OpenStudy (anonymous):

combine like terms and put terms in order of the exponent of x starting with the biggest exponent. So the term with x is first and the constant is last: f(x)=5x-4

OpenStudy (anonymous):

(b) Find the coordinates of the vertex and indicate whether it corresponds to the maximum or the minimum of f. (c) Find the x- and y-intercepts.

OpenStudy (anonymous):

oh and the equation is x squared not x 2

OpenStudy (anonymous):

Standard form is Ax² + Bx + C = 0

OpenStudy (anonymous):

@jonzion1122 Hellooo!

OpenStudy (anonymous):

Ok - More? to find the vertex use the following formula x = (-b/(2a). b=4, a=1 from the quadratic so x=-2, sub -2 into the quadratic to get your value of y: y=(-2)^2+4(-2)-6 to find the y=intercept (no matter what function you have) substitute 0 for x and do the math to find x-intercepts, sub y to 0 and solve for x. In a quadratic function you can always use the quadratic formula or completing the square and that's the case for this function.